A compound is made of copper, nitrogen, and oxygen atoms. A sample of the compound has 6.4 grams of copper, 2.8 grams of nitrogen and 9.6 grams of oxygen. The compound has a molar mass of 188 g /mol. Write the name and formula of the compound.
1
Expert's answer
2017-03-13T15:59:06-0400
M (Cu) = 63.6 g/mol

M (N) = 12 g/mol

M (O) = 16 g/mol

n (Cu) = 6.4 / 63.6 = 0.1 mol

n (N) = 2.8 / 12 = 0.23 mol

n (O) = 9.6 / 16 = 0.6 mol

Formula: CuN 2 O 6 – Copper Dinitrate